OpenStudy (anonymous):

You buy clothing at a sale you buy sneakers a 2/3 of their original price of $60, a sweater a 7/10 of its original price of $35, and a jacket at 5/6 of its original price of $45. How much money did you spend? how much money did you save? what fraction of the original price di you save? @bohotness @chosenmatt I will give 4 medals.

OpenStudy (bohotness):

http://www.jiskha.com/display.cgi?id=1359336968

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{3}\ of \ \$60\implies \cfrac{2}{3}\cdot 60\implies \square ? \\ \quad \\ \cfrac{7}{10}\ of \ \$35\implies \cfrac{7}{10}\cdot 35\implies \square ? \\ \quad \\ \cfrac{5}{6}\ of \ \$45\implies \cfrac{5}{6}\cdot 45\implies \square ? \\ \quad \\ thus\qquad \square ?+ \square ?+ \square ?=?\)